# Avat3r: Large Animatable Gaussian Reconstruction Model for High-fidelity 3D Head Avatars

### v1

@@ Abstract
Traditionally, creating photo-realistic 3D head avatars requires a studio-level multi-view capture setup and expensive optimization during test-time, limiting the use of digital human doubles to the VFX industry or offline renderings. To address this shortcoming, we present Avat3r, which regresses a high-quality and animatable 3D head avatar from just a few input images, vastly reducing compute requirements during inference. More specifically, we make Large Reconstruction Models animatable and learn a powerful prior over 3D human heads from a large multi-view video dataset. For better 3D head reconstructions, we employ position maps from DUSt3R and generalized feature maps from the human foundation model Sapiens. To animate the 3D head, our key discovery is that simple cross-attention to an expression code is already sufficient. Finally, we increase robustness by feeding input images with different expressions to our model during training, enabling the reconstruction of 3D head avatars from inconsistent inputs, e.g., an imperfect phone capture with accidental movement, or frames from a monocular video. We compare Avat3r with current state-of-the-art methods for few-input and single-input scenarios, and find that our method has a competetive advantage in both tasks. Finally, we demonstrate the wide applicability of our proposed model, creating 3D head avatars from images of different sources, smartphone captures, single images, and even out-of-domain inputs like antique busts.
@@ 1 Introduction
Lifelike 3D head avatars find use in a variety of applications such as telepresence, movie production, personalized games, or digital education. Typically, the creation of a high-quality avatar requires studio-quality videos that have to be recorded in hour-long sessions in expensive capture setups. This limits the applicability of a 3D head avatar creation pipeline, especially in casual settings where users may only want to use a few images taken with their smartphone to create a convincing digital double. In this paper, we present Avat3r, a method that can reconstruct an animatable 3D head avatar from just a few images of a person.
@@
Building a photo-realistic 3D head avatar from a few casual images involves several hard challenges. First, sparse 3D reconstruction poses a heavily under-constrained optimization problem since some regions such as the mouth interior or the side of the head may have simply not been seen in the available images. Next, animating the head requires plausibly morphing the reconstructed face without ever having seen that particular person talking or sticking their tongue out. Finally, the input images may be inconsistent, e.g., because the person could not remain completely static during a casual capture. All of these three challenges, i.e., sparse 3D reconstruction, face animation, and robust reconstruction, have to be solved simultaneously to successfully reconstruct a photo-realistic 3D head avatar given just a few images.
@@
Due to the aforementioned challenges, the current best methods for high-quality avatar creation [16, 39, 37, 26] cannot be applied since their optimization-based approach requires many more observations than just a few images of a person. On the other hand, methods that aim at 3D-aware portrait animation [14, 43, 7, 8] learn a prior over 3D heads from large video datasets that include a variety of different people. As a result, existing methods can already convincingly animate faces from as few as just a single input image. However, such methods typically focus on frontal facial appearance and may sacrifice 3D-consistency or temporal consistency to reach satisfying image synthesis quality. This makes these methods less applicable in scenarios where a proper 3D model is required. Another approach is to build a photo-realistic 3D head model that learns the distribution of 3D face geometry, expression, and appearance, akin to classic 3DMMs. While such unconditional generative models are very versatile and recent works have shown promising results [50, 58], they typically suffer from limited generalizability along the identity axis due to existing 3D face datasets only providing a few hundred different persons [25, 34, 51, 36].
@@
Our approach is motivated by the observation that 3D data for faces is limited in the identity axis, but not in the expression axis that is already well captured in existing datasets. We therefore design a system that is conditional on the identity of a person, but generalizes along the expression axis. Opposed to photo-realistic 3D face models, the network does not have to learn the full extent of human facial appearance but just needs to understand how to reconstruct it from a few example images, which simplifies the task. Inspired by recent Large Reconstruction Models [40, 49, 56], we devise an architecture that predicts 3D Gaussians for each pixel in the input images. As such, we deliberately avoid anchoring the 3D Gaussians on a template mesh, as common in the 3D Head Avatar domain [37, 39, 3]. The reason is that human facial appearance is complex and diverse. Limiting the 3D representation to a fixed number of Gaussians anchored on a template mesh with uniform topology will cause issues in regions such as hair or face accessories. Instead, when predicting Gaussians for each foreground pixel, a person with voluminous hair will receive more primitives to model their head than a bald person. To further simplify the task of sparse 3D reconstruction, we first compute position maps for each input image via DUSt3R [46]. These position maps are injected into the network via skip connections and act as a coarse starting position for each Gaussian. We surprisingly find that DUSt3R is still capable of producing reasonable position maps even if the input images are not consistent, e.g., because they show different facial expressions or are the result of a view-inconsistent image generation model. We exploit this remarkable capability of DUSt3R to train our model on input images taken at different timesteps, which not only enables training and inference on monocular video datasets but also makes the model more robust to inconsistencies in the input images. Such inconsistencies can quickly occur in a real-life scenario where a person may not stay completely still during a phone capture of their head. Finally, to model the animation of the face, we find that simple cross-attention between intermediate feature maps and a descriptive expression code are already sufficient to generalize over the space of facial expressions. Taken together, our contributions are as follows:
@@
  • We build the first animatable large 3D reconstruction model and show that simple cross-attention on expression token sequences is sufficient to model complex facial animations.
  • To improve generalization and robustness, we integrate priors from Dust3r and Sapiens via skip connections, and feed input images from different timesteps to the network during training.
@@ 2.1 Large 3D Reconstruction Models
Recently, generalized feed-forward models for 3D reconstruction from sparse input views have garnered considerable attention due to their applicability in heavily under-constrained scenarios. The Large Reconstruction Model (LRM) [18] uses a transformer-based encoder-decoder pipeline to infer a NeRF reconstruction from just a single image. Newer iterations have shifted the focus towards generating 3D Gaussian representations from four input images [40, 49, 56, 5, 6, 31], showing remarkable novel view synthesis results. The paradigm of transformer-based sparse 3D reconstruction has also successfully been applied to lifting monocular videos to 4D [38].
Yet, none of the existing works in the domain have studied the use-case of inferring animatable 3D representations from sparse input images, which is the focus of our work. To this end, we build on top of the Large Gaussian Reconstruction Model (GRM) [49].
@@
A different line of work focuses on animating portraits in a 3D-aware manner. MegaPortraits [14] builds a 3D Volume given a source and driving image, and renders the animated source actor via orthographic projection with subsequent 2D neural rendering. 3D morphable models (3DMMs) [2] are extensively used to obtain more interpretable control over the portrait animation. For example, StyleRig [41] demonstrates how a 3DMM can be used to control the data generated from a pre-trained StyleGAN [20] network. ROME [22] predicts vertex offsets and texture of a FLAME [29] mesh from the input image. A TriPlane representation is inferred and animated via FLAME [29] in multiple methods like Portrait4D [11], Portrait4D-v2 [12], and GPAvatar [8]. Others, such as VOODOO 3D [43] and VOODOO XP [42], learn their own expression encoder to drive the source person in a more detailed manner.
All of the aforementioned methods require nothing more than a single image of a person to animate it. This allows them to train on large monocular video datasets to infer a very generic motion prior that even translates to paintings or cartoon characters. However, due to their task formulation, these methods mostly focus on image synthesis from a frontal camera, often trading 3D consistency for better image quality by using 2D screen-space neural renderers. In contrast, our work aims to produce a truthful and complete 3D avatar representation from the input images that can be viewed from any angle.
@@ 2.3 Photo-realistic 3D Face Models
The increasing availability of large-scale multi-view face datasets [25, 34, 36, 51] has enabled building photo-realistic 3D face models that learn a detailed prior over both geometry and appearance of human faces. HeadNeRF [17] conditions a Neural Radiance Field (NeRF) [35] on identity, expression, albedo, and illumination codes. VRMM [52] builds a high-quality and relightable 3D face model using volumetric primitives [32]. One2Avatar [54] extends a 3DMM by anchoring a radiance field to its surface. More recently, GPHM [50] and HeadGAP [58] have adopted 3D Gaussians to build a photo-realistic 3D face model.
Photo-realistic 3D face models learn a powerful prior over human facial appearance and geometry, which can be fitted to a single or multiple images of a person, effectively inferring a 3D head avatar. However, the fitting procedure itself is non-trivial and often requires expensive test-time optimization, impeding casual use-cases on consumer-grade devices. While this limitation may be circumvented by learning a generalized encoder that maps images into the 3D face model’s latent space, another fundamental limitation remains. Even with more multi-view face datasets being published, the number of available training subjects rarely exceeds the thousands, making it hard to truly learn the full distibution of human facial appearance. Instead, our approach avoids generalizing over the identity axis by conditioning on some images of a person, and only generalizes over the expression axis for which plenty of data is available.
@@ 3 Method
where $I=\{I_{1},...,I_{V}\}$ are the $V$ input images, $\pi$ their corresponding camera parameters, and $z_{exp}$ a code describing the desired facial expression. In our experiments, $V=4$. Note that the images $I$ do not necessarily have to stem from multi-view camera systems or synthetic multi-view renderings of objects as it is common for large reconstruction models. Instead, the input images may show the scene at different instances in time, making the model amenable for training and inference on frames from monocular videos.
@@ 4.1 Training
Avat3r is trained with Adam [24] using a learning rate of $55e-$ 5. We only introduce $\mathcal{L}_{lpips}$ after 3M optimization steps to avoid focusing on high-frequency details too early. In total, the model is trained for 3.5M steps with a batch size of 1 per GPU on 8 A100 GPUs which takes roughly 4 days. We further find that training convergence is significantly improved when supervising multiple viewpoints for each batch. I.e., a batch consists of 12 images: 4 with random expression that are used as input, and 8 viewpoints from the target expression that are used as supervision.
@@ 4.2 Experiment Setup
We evaluate the ability to create a 3D head avatar from one or multiple images of an unseen person.
@@
We employ three paired-image metrics to measure the quality of individual rendered images: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [47], and Learned Perceptual Image Patch Similarity (LPIPS) [57]. Furthermore, we measure visual similarity of a rendered video to its ground truth counterpart with the perceptual video metric JOD [33]. Finally, we make use of two face-specific metrics: Average Keypoint Distance (AKD) measured in pixels with keypoints estimated from PIPNet [19], and cosine similarity (CSIM) of identity embeddings based on ArcFace [10].
@@
We compare our method with recent state-of-the-art systems for 3D head avatar creation from one and few input images.
@@
GAGAvatar [7]. A method for 3D-aware portrait animation from a single image that constructs two sets of 3D Gaussians: The first set is obtained by predicting 2 Gaussians for each pixel in the input image with a CNN, the second set is decoded from FLAME vertices and automatically animated through the FLAME mesh.
Image Video Face
@@
4-in GPAvatar 20.3 0.703 0.304 4.62 6.08 0.277
Ours 22.0 0.732 0.309 5.20 5.93 0.595
Table 1: 3D Head Avatar Creation Comparison. Our method outperforms GPAvatar [8] in almost all metrics.
Input images GPAvatar [8] Ours GT
Figure 4: Comparison for few-shot 3D Head Avatar creation. We compare the ability of Avat3r to create 3D head avatars from four input images against GPAvatar [8]. Our method generalizes better and produces higher quality avatars.
4.3 Few-shot 3D Head Avatar Creation
We compare the ability to create a realistic 3D head avatar from just four input images with GPAvatar [8]. To this end, we train GPAvatar on the Ava256 dataset in the same setting as our model and use the high-quality tracked meshes provided with the dataset to drive GPAvatar. We evaluate both methods at $512\times 512$ pixels resolution on the same set of 12 hold-out persons, sampling in total 1000 different target expressions from different viewpoints. For the video-based metric JOD, we render one video of 100 frames from the frontal camera for each hold-out person. The quantitative results are listed in Tab. 1. It can be seen that our method clearly outperforms the baseline. We observe the biggest difference in metrics measuring alignment (PSNR, SSIM), temporal consistency (JOD), and identity similarity (CSIM). This indicates that Avat3r produces 3D head avatars that better resemble the person in the input images and produces smoother video renderings since it does not require a screen-space super-resolution network as GPAvatar. We further show qualitative comparisons in Fig. 4. Our method produces more expressive avatars with higher rendering quality than the baseline and even performs well on a person with colored hair who wears a headband.
To further assess Avat3r’s generalization capabilities, we create few-shot 3D head avatars for subjects from the multi-view video dataset NeRSemble [25] which has not been used during training. The results can be seen in Fig. 3. Note that the images in the NeRSemble dataset exhibit entirely different characteristics in terms of lighting, camera parameters, and viewpoints, compared to the training images from the Ava256 dataset. Nevertheless, Avat3r generates realistic 3D head avatars in this setting as well.
Input GAGavatar GAGavatar Ours GT
Figure 5: Comparison for single-shot 3D Head Avatar creation. We compare Avat3r with the recent 3D-aware portrait animation methods GAGAvatar [7] in a self-reenactment scenario on hold-out persons from the Ava256 dataset. GAGavatar denotes a version of the baseline that we trained on the Ava256 dataset. Our method shows better rendering quality especially for extreme expressions and viewing angles.
4.4 Single-shot 3D Head Avatar Creation
To make Avat3r amenable for inference on only a single input image, we make use of a pre-trained 3D GAN [55] to first lift the single image to 3D and then render four views of the head. These renderings then constitute the input for Avat3r. We conduct comparisons with the recent 3D-aware portrait animation method GAGAvatar [7]. Specifically, we compare with two version of GAGAvatar: One provided by the authors which is trained on VFHQ [48], and another version, denoted as GAGAvatar, that we train on the Ava256 dataset in the same setting as our method. To drive GAGAvatar, we use their monocular FLAME tracker to obtain tracked meshes. Fig. 5 shows qualitative comparisons between our method and GAGAvatar for single input images of hold out persons. Note that our method performs competitively compared to the single-input baselines despite never being trained for a single-shot scenario. We find that Avat3r produces more realistic facial expressions than GAGAvatar which is limited by FLAME’s expression space. Furthermore, our method allows much more extreme viewpoint changes without sacrificing rendering quality.

### v2

@@ Abstract
Traditionally, creating photo-realistic 3D head avatars requires a studio-level multi-view capture setup and expensive optimization during test-time, limiting the use of digital human doubles to the VFX industry or offline renderings. To address this shortcoming, we present Avat3r, which regresses a high-quality and animatable 3D head avatar from just a few input images, vastly reducing compute requirements during inference. More specifically, we make Large Reconstruction Models animatable and learn a powerful prior over 3D human heads from a large multi-view video dataset. For better 3D head reconstructions, we employ position maps from DUSt3R and generalized feature maps from the human foundation model Sapiens. To animate the 3D head, our key discovery is that simple cross-attention to an expression code is already sufficient. Finally, we increase robustness by feeding input images with different expressions to our model during training, enabling the reconstruction of 3D head avatars from inconsistent inputs, e.g., an imperfect phone capture with accidental movement, or frames from a monocular video. We compare Avat3r with current state-of-the-art methods for few-input and single-input scenarios, and find that our method has a competitive advantage in both tasks. Finally, we demonstrate the wide applicability of our proposed model, creating 3D head avatars from images of different sources, smartphone captures, single images, and even out-of-domain inputs like antique busts.
@@ 1 Introduction
Lifelike 3D head avatars find use in a variety of applications such as telepresence, movie production, personalized games, or digital education. Typically, the creation of a high-fidelity avatar requires studio-quality videos that have to be recorded in hour-long sessions in expensive capture setups. This limits the applicability of a 3D head avatar creation pipeline, especially in casual settings where users may only want to use a few images taken with their smartphone to create a convincing digital double. In this paper, we present Avat3r, a method that can reconstruct an animatable 3D head avatar from just a few images of a person.
@@
Building such a system involves several hard challenges. First, sparse 3D reconstruction poses a heavily under-constrained optimization problem since some regions such as the mouth interior or the side of the head may have simply not been seen in the available images. Next, animating the head requires plausibly morphing the reconstructed face without ever having seen that particular person talking or sticking their tongue out. Finally, the input images may be inconsistent, e.g., because the person could not remain completely static during a casual capture. All of these three challenges, i.e., sparse 3D reconstruction, face animation, and robust reconstruction, have to be solved simultaneously to successfully reconstruct a photo-realistic 3D head avatar given just a few images.
@@
Due to the aforementioned challenges, the current best methods for high-quality avatar creation [23, 51, 49, 36, 58, 59, 30, 85, 70] cannot be applied since their optimization-based approach requires many more observations than just a few images of a person. On the other hand, methods that aim at 3D-aware portrait animation [20, 62, 13, 14] learn a prior over 3D heads from large video datasets. As a result, existing methods can already convincingly animate faces from as few as just a single input image. However, such methods typically focus on frontal facial appearance and may sacrifice 3D-consistency or temporal consistency to reach satisfying image synthesis quality. This makes these methods less applicable in scenarios where a proper 3D model is required. Another approach is to build a photo-realistic 3D head model that learns the distribution of 3D face geometry, expression, and appearance, akin to classic 3DMMs. While such unconditional generative models are very versatile and recent works have shown promising results [72, 81], they typically suffer from limited generalizability along the identity axis due to existing 3D face datasets only providing a few hundred different persons [35, 46, 73, 48].
@@
Our approach is motivated by the observation that 3D data for faces is limited in the identity axis, but not in the expression axis that is already well captured in existing datasets. We therefore design a system that is conditional on the identity of a person, but generalizes along the expression axis. Opposed to photo-realistic 3D face models, the network does not have to learn the full extent of human facial appearance but just needs to understand how to reconstruct it from a few example images, which simplifies the task. Inspired by recent Large Reconstruction Models [56, 71, 78], we devise an architecture that predicts 3D Gaussians for each pixel in the input images. As such, we deliberately avoid anchoring the 3D Gaussians on a template mesh, as common in the 3D Head Avatar domain [49, 51, 6]. The reason is that human facial appearance is complex and diverse. Limiting the 3D representation to a fixed number of Gaussians anchored on a template mesh with uniform topology will cause issues in regions such as hair or face accessories. Instead, when predicting Gaussians for each foreground pixel, a person with voluminous hair will receive more primitives to model their head than a bald person. To simplify the task of sparse 3D reconstruction, we enrich each input image with feature maps from Sapiens [33] and position maps from DUSt3R [65]. These position maps are injected into the network via skip connections and act as a coarse starting position for each Gaussian. We find that DUSt3R is still capable of producing reasonable position maps even if the input images are not consistent. We exploit this remarkable capability of DUSt3R to train our model on input images taken at different timesteps, which not only enables training and inference on larger monocular video datasets but also makes the model more robust to inconsistencies in the input images. Such inconsistencies can quickly occur in a real-life scenario where a person may not stay completely still during a phone capture of their head. Finally, to model the animation of the face, we find that simple cross-attention between intermediate feature maps and a descriptive expression code are already sufficient to generalize over the space of facial expressions. Taken together, our contributions are as follows:
@@
  • We build the first animatable large 3D reconstruction model and show that simple cross-attention on expression codes is sufficient to model complex facial animations.
  • To improve robustness, we integrate priors from Dust3r and Sapiens via skip connections, and feed input images from different timesteps to the network during training.
@@ 2.1 Large 3D Reconstruction Models
Recently, generalized feed-forward models for 3D reconstruction from sparse input views have garnered considerable attention due to their applicability in heavily under-constrained scenarios. The Large Reconstruction Model (LRM) [28] uses a transformer-based encoder-decoder pipeline to infer a NeRF reconstruction from just a single image. Newer iterations have shifted the focus towards generating 3D Gaussian representations from four input images [56, 71, 78, 10, 12, 42], showing remarkable novel view synthesis results. Very recently, FaceLift [45] has proposed a combination of personalized multi-view image diffusion and a large reconstruction model for single-image 3D head reconstruction. The paradigm of transformer-based sparse 3D reconstruction has also successfully been applied to lifting monocular videos to 4D [50].
Yet, none of the existing works in the domain have studied the use-case of inferring animatable 3D representations from sparse input images, which is the focus of our work.
2.2 Monocular 3D Head Avatar Reconstruction
A popular task is to reconstruct an animatable 3D head avatar from a monocular input video. Several methods have been proposed based on meshes [24], NeRFs [22, 86, 69], SDFs [82], points [83], and Gaussians [67, 11]. However, due to their optimization-based nature, these approaches often exhibit poor viewpoint extrapolation capabilities due to overfitting on the input viewpoint. Very recently,  Tang et al. [55] proposed to improve the geometric fidelity of reconstructed monocular avatars by incorporating priors from a personalized image diffusion model, still requiring a full monocular input video. In contrast, our method requires only 4 input images to reconstruct a complete and animatable 3D head with extensive viewpoint coverage.
@@ 2.2 Monocular 3D Head Avatar Reconstruction
A different line of work focuses on animating portraits in a 3D-aware manner, based on 3D volumes [20], neural rendering [60], meshes [32], TriPlanes [17, 18, 14, 62, 61], or 3D Gaussians [13]. Typically, the animation control is injected via a 3D morphable model (3DMM) [4, 40]. Concurrent to our work, CAP4D [57] uses personalized multi-view image diffusion and LAM [26] uses cross attention to build a 3D head avatar from a single image. While the aforementioned methods can animate a single portrait, they mostly focus on image synthesis from a frontal camera, often trading 3D consistency for better image quality by using 2D screen-space neural renderers. In contrast, our work aims to produce a truthful and complete 3D avatar representation from the input images that can be viewed from any angle.
@@ 2.4 Photo-realistic 3D Face Models
The increasing availability of large-scale multi-view face datasets [35, 46, 48, 73, 5] has enabled building photo-realistic 3D face models that learn a detailed prior over both geometry and appearance of human faces. HeadNeRF [27] conditions a Neural Radiance Field (NeRF) [47] on identity, expression, albedo, and illumination codes. VRMM [74] builds a high-quality and relightable 3D face model using volumetric primitives [44]. One2Avatar [76] extends a 3DMM by anchoring a radiance field to its surface. More recently, GPHM [72] and HeadGAP [81] have adopted 3D Gaussians to build a photo-realistic 3D face model. Another line of work utilizes 3D Generative Adversarial Networks [7, 8, 25] to learn the distribution of animatable head avatars [53, 54, 38, 2] from frontal 2D image datasets.
To obtain the final 3D avatar, typically a costly fitting process is necessary, impeding casual use-cases on consumer-grade devices. While this limitation may be circumvented by learning a generalized encoder that maps images into the 3D face model’s latent space [80, 52, 43, 77, 3], another fundamental limitation remains. Even with more multi-view face datasets being published, the number of available training subjects rarely exceeds the thousands, making it hard to truly learn the full distribution of human facial appearance. 3DGAN-based approaches mitigate this by training on single image datasets, but typically struggle with view extrapolation. In this paper, we pursue a different approach that avoids generalizing over the identity axis, by conditioning on some images of a person, and only generalizes over the expression axis, for which plenty of data is available.
@@ 3 Method
where $I=\{I_{1},...,I_{V}\}$ are the $V$ input images, $\pi$ their corresponding camera parameters, and $z_{exp}$ a code describing the desired facial expression, such as an expression code of FLAME [40]. In our experiments, $V=4$. Note that the images $I$ do not necessarily have to stem from multi-view camera systems or synthetic multi-view renderings of objects as it is common for large reconstruction models. Instead, the input images may show the scene at different instances in time, making the model amenable for training and inference on frames from monocular videos.
@@ 4.1 Training
Avat3r is trained with Adam [34] using a learning rate of $5\text{e-}5$. We only introduce $\mathcal{L}_{lpips}$ after 3M optimization steps to avoid focusing on high-frequency details too early. In total, the model is trained for 3.5M steps with a batch size of 1 per GPU on 8 A100 GPUs which takes roughly 4 days. We further find that training convergence is significantly improved when supervising multiple viewpoints for each batch. I.e., a batch consists of 12 images: 4 with random expression that are used as input, and 8 viewpoints from the target expression that are used as supervision.
@@ 4.2 Experiment Setup
We evaluate the ability to create a 3D head avatar from four images of an unseen person.
@@
We employ three paired-image metrics to measure the quality of individual rendered images: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [66], and Learned Perceptual Image Patch Similarity (LPIPS) [79]. Furthermore, we make use of two face-specific metrics: Average Keypoint Distance (AKD) measured in pixels with keypoints estimated from PIPNet [29], and cosine similarity (CSIM) of identity embeddings based on ArcFace [16].
@@
We compare our method with recent state-of-the-art systems for 3D head avatar creation from few input images.
HeadNeRF [27]. A NeRF-based Autodecoder with disentangled identity and expression spaces that can be fitted to an arbitrary number of images.
InvertAvatar [80]. A feed-forward encoder that predicts the underlying representation of Next3D [53] from 4 input images which then can be animated.
@@
$\overbrace{\hskip 150.79968pt}^{\text{\normalsize Ava256 dataset}}$
$\overbrace{\hskip 150.79968pt}^{\text{\normalsize NeRSemble dataset}}$
Input images HeadNeRF [27] InvertAvatar [80] GPAvatar [14] Ours GT
Figure 3: Comparison for few-shot 3D Head Avatar creation. We evaluate the ability of Avat3r to create 3D head avatars from four input images in a self-reenactment setting. Note that the NeRSemble dataset has not been used during training of Avat3r and therefore both source and driver person are out-of-domain in this case. Nevertheless, our method produces high quality avatars.
4.3 Few-shot 3D Head Avatar Creation
We compare the ability to create a realistic 3D head avatar from just four input images on both the Ava256 dataset [46] and the NeRSemble dataset [35]. To this end, we train GPAvatar on the Ava256 dataset in the same setting as our model and use the high-quality tracked meshes provided with the dataset to drive GPAvatar. For HeadNeRF and InvertAvatar, we use the published code bases for reconstruction, animation and rendering. We evaluate all methods at $512\times 512$ pixels resolution on the same hold-out persons. For evaluation on Ava256, we sample 1000 different target expressions from 12 persons from different viewpoints. On NeRSemble, we sample 100 target images from 3 male and 3 female subjects.
The quantitative results are listed in Tab. 1. It can be seen that our method clearly outperforms the baselines. We observe large differences in metrics measuring alignment (PSNR, SSIM), identity similarity (CSIM), and facial keypoint similarity (AKD). This indicates that Avat3r produces 3D head avatars that better resemble the person in the input images and produces more realistic animations. Note that HeadNeRF cannot render any other background than white. We therefore compute metrics against ground truth images with white background for that method. We further show qualitative comparisons in Fig. 3. Our method produces more expressive avatars with higher rendering quality than the baselines and even performs well on a person with colored hair who wears a headband. Also, note that the NeRSemble dataset has not been used during training and the images in this dataset exhibit entirely different characteristics in terms of lighting, camera parameters, and viewpoints, compared to the training images from the Ava256 dataset. Nevertheless, Avat3r generates realistic 3D head avatars in this setting as well.
For video renderings and single-image results, see our supplemental material.
@@ 4.3 Few-shot 3D Head Avatar Creation
Ava256
HeadNeRF 9.1 0.64 0.52 6.9 0.11
InvertAvatar 14.2 0.36 0.55 15.8 0.29
GPAvatar 19.4 0.69 0.34 5.3 0.31
Ours 20.7 0.71 0.33 4.8 0.59
NeRSemble
HeadNeRF 9.8 0.69 0.47 4.9 0.22
InvertAvatar 17.2 0.44 0.49 5.6 0.44
GPAvatar 17.6 0.67 0.40 5.7 0.07
Ours 20.5 0.75 0.33 3.7 0.50
Table 1: 3D Head Avatar Creation Comparison. Our method outperforms the baselines in all metrics.